OpenStudy (anonymous):

Evaluate the expression when a = 4. 2a^3 A. 14 B. 24 C. 128 D. 512

OpenStudy (anonymous):

128. 2*4^3=2*4*4*4=2*64=128

OpenStudy (richyw):

oops, typo there